OpenStudy (starryknight):

Solve for x in the following system of equations. -6x - 5y = 3 7x + 9y = -32 -18 7 -36 2 please help me find y please and thankyou :))))

OpenStudy (anonymous):

Umm. I am getting that x is 7 and y is -9. Are you sure you are looking for y? 9 isnt one of the options.

OpenStudy (starryknight):

oops your right sorry im looking for x sorry :))))

OpenStudy (anonymous):

-18 7 -36 2

OpenStudy (anonymous):

waht do you mean, bu this

OpenStudy (anonymous):

x is 7.

OpenStudy (nirmalnema):

-6x - 5y = 3 7x + 9y = -32 x=-5y-3 / 6 x=-32-9y /7 -5y-3 / 6 = -32-9y /7 (5y+3)7=6(32+9y) 35y+21=192+54y 54y-35y=21-192 -171=19y y=9 put this value in any equation you will get x = 7

OpenStudy (starryknight):

-18 7 -36 2 these are the answer choices

OpenStudy (starryknight):

thank you

OpenStudy (nirmalnema):

you welcome//!!!

OpenStudy (anonymous):

ok it's solved